# Quantum-Classical Hybrid Quantized Neural Network

### v2

@@ Appendix A Error Bound of Piecewise Linear Approximation in FNN
To determine $n$, assume $\sigma$ has $\|\sigma^{\prime\prime}\|_{\infty}\leq M$. Construct $\hat{\sigma}$ over $[-R,R]$ with $n$ equal segments, each of width $h=\frac{2R}{n}$. The interpolation error is:
@@
Appendix B Spline Quantumization Protocol
Quantum computing faces the formidable challenge of not only optimizing QUBO models but also addressing highly nonlinear optimization problems as optimization targets. The significance of this issue lies in its ubiquity across diverse scientific and industrial domains, where real-world problems often exhibit intricate, arbitrary relationships among variables. This fact underscores the urgency for quantum algorithms capable of navigating and optimizing complex, highly nonlinear landscapes.
Formally we consider the following problem:
@@ Appendix B Spline Quantumization Protocol
where $f(\cdot)$ is a function of arbitrary form. This formal definition of the optimization problem unlockes the further potential of quantum computing in solving real-world optimization problems. A powerful approach to approximate highly nonlinear functions is to use spline interpolation. Splines are piecewise linear functions that provide a smooth and flexible fit to data points. Spline interpolation can transform the optimization landscape into a more tractable form, making it amenable to quantum optimization algorithms. The spline fitting process involves the following steps:
@@
B.1 Piecewise Constant Segment
@@
We have now derived a QUBO model capable of approximately solving optimization problems with arbitrary objective functions in (B). However, it is important to note that in representing $\beta_{i}$, we introduced two inequality constraints. Here, we introduce the following two approaches:
@@
    It is important to note that the higher order terms cancel out when the expression is expanded.
@@
Building upon the versatility of spline approximation in simplifying complex optimization landscapes, we use the following sign function as an example to illustrates the efficiency of our approach.
@@
$\text{sign}(x)$ ($1$ if $x\geq 0$ else $-1$) can be represented as
@@
Compared with the model in [54], we do not need to introduce the intermediate variable. Additionally, when representing the sign activation function, we do not need to introduce auxiliary variables for order reduction, as done in [54]. These optimizations can save $O(WDN\log W)$ bits for us, where $W,D,N$ denote the network width, network depth and dataset size respectively. Indeed we are able to reduce the coefficients of the highest-order terms in the model’s qubit count by 50%.
@@
Appendix C Proof of Convergence of QCGD with Random Quantum Oracle
@@
Proof: Similar as [45], the algorithm utilizes an augmented Lagrangian function, defined as
@@
$\mathcal{Q}_{\alpha}(\mathbf{V},\mathbf{z})=\mathrm{Tr}(\mathbf{C}\mathbf{V})+\mathbf{z}^{\top}(\mathcal{L}\mathbf{V}-\mathbf{v})+\frac{\alpha}{2}\|\mathcal{L}\mathbf{V}-\mathbf{v}\|^{2}\quad\text{for}\leavevmode\nobreak\ \mathbf{V}\in\Delta^{p},$ (65)
@@
The QCGD algorithm begins by initializing $\alpha$, $\mathbf{V}$ and $\mathbf{z}$, and then iterates through primal and dual updates. In the primal step, the algorithm computes the gradient of the augmented Lagrangian with respect to $\mathbf{V}$ and finds a direction that minimizes the linearized loss. This step involves solving a standard QUBO problem to determine the update direction. The primal variable $\mathbf{V}$ is then updated by taking a step towards this direction. In the dual step, the gradient of the augmented Lagrangian with respect to $\mathbf{z}$ is computed, and $\mathbf{z}$ is updated using gradient ascent. Over iterations, the penalty parameter $\alpha$ is increased as $\alpha_{i}=\alpha_{0}\sqrt{\delta i+1}$ to ensure that $\mathbf{V}$ converges to a feasible solution that satisfies the constraints.
@@
[…]
@@
Appendix D Analysis of Total Error Due to Truncation
The energy function of the Ising model is given by:
$E[\sigma|J_{ij},h_{i}]=-\sum_{i<j}J_{ij}\sigma_{i}\sigma_{j}-\sum_{i}h_{i}\sigma_{i}.$ (86)
The objective function of QUBO model is
$f(\mathbf{x}|q_{ij})=\sum_{i\neq j}q_{ij}x_{i}x_{j}+\sum_{i}q_{i}x_{i}.$ (87)
Let $\sigma^{*}$ be the optimal configurations under origin Ising coefficients $J_{ij}$s and $h_{i}$s, $\mathbf{x}^{*}$ be the corresponding optimal solution to the QUBO model with origin QUBO coefficients $q_{ij}$s and $q_{i}$s. We use ${\sigma^{*}}^{\prime}$ to denote the optimal configurations under truncated Ising coefficients, and let ${\mathbf{x}^{*}}^{\prime}$ be the corresponding solution in QUBO model. Then we have the following conclusion
Lemma 3
If the Ising coefficients are truncated to $d$ bits, the error due to truncation is in the order of
$\Delta\text{Hamiltonian}$ $=|E[\sigma|J_{ij},h_{i}]-E[{\sigma^{*}}^{\prime}|J_{ij},h_{i}]|=O(2^{-d}\cdot n^{2}),$ (88)
ΔQUBO $\Delta\text{QUBO Value}$ $=|f(\mathbf{x}|q_{ij})-f({\mathbf{x}^{*}}^{\prime}|q_{ij})|=O(2^{-d}\cdot n^{2}).$ (89)
Proof: The change in the Ising coefficients due to truncations is
$|J_{ij}-J^{\prime}_{ij}|,|h_{i}-h^{\prime}_{i}|=O(2^{-d}),$ (90)
consequently for any $\sigma$
$|E[\sigma|J_{ij},h_{i}]-E[\sigma|J^{\prime}_{ij},h^{\prime}_{i}]|=O(2^{-d}\cdot n^{2}).$ (91)
Note that
$E[{\sigma^{*}}^{\prime}|J_{ij},h_{i}]\leq$ $E[{\sigma^{*}}^{\prime}|J^{\prime}_{ij},h^{\prime}_{i}]+O(2^{-d}\cdot n^{2})$
$\leq$ $E[\sigma^{*}|J^{\prime}_{ij},h^{\prime}_{i}]+O(2^{-d}\cdot n^{2}),$ (92)
and
$E[{\sigma^{*}}^{\prime}|J_{ij},h_{i}]\geq$ $E[\sigma^{*}|J_{ij},h_{i}]$
$\geq$ $E[\sigma^{*}|J^{\prime}_{ij},h^{\prime}_{i}]-O(2^{-d}\cdot n^{2}),$ (93)
we can obtain (88) by combining (D) and (D).
To convert the Ising model to the QUBO model, we replace the spin variables $\sigma_{i}$ with binary variables $x_{i}$, by using the following equation:
$\sigma_{i}=2x_{i}-1$ (94)
Substitute this into the Ising model energy function:
$E$ $=-\sum_{i<j}J_{ij}(2x_{i}-1)(2x_{j}-1)-\sum_{i}h_{i}(2x_{i}-1)$ (95)
$=-\sum_{i<j}J_{ij}\left[4x_{i}x_{j}-2x_{i}-2x_{j}+1\right]-\sum_{i}h_{i}\left[2x_{i}-1\right]$ (96)
$=-\sum_{i<j}4J_{ij}x_{i}x_{j}+\sum_{i}(2\sum_{j\neq i}J_{ij}-2h_{i})x_{i}-\sum_{i<j}J_{ij}+\sum_{i}h_{i}.$ (97)
Hence $E[\sigma|J_{ij},h_{i}]=f(\mathbf{x}|q_{ij})+(\sum_{i}h_{i}-\sum_{i<j}J_{ij})$ and
ΔQUBO $\Delta\text{QUBO Value}=\Delta\text{Hamiltonian}+O(2^{-d}\cdot n^{2})=O(2^{-d}\cdot n^{2}).$ (98)
When $d=\Omega(\log n)$, the additive error is in the order of $O(\frac{1}{\sqrt{t}})$ for all $t\leq T$, and thus the convergence follows from Proposition 1. $\Box$
Appendix E Additional Details of Experimental Results
@@
The training set consists of $12000$ images ($6000$ coat and $6,000$ sandal images), and the test set consists of $2000$ images ($1000$ coat and $1000$ sandal images). Each image is divided into $3$ columns vertically, resulting in two $28\times 9$ grid and one $28\times 10$ grid. We then calculate the number of zero pixels in each column and use two threshold values to determine the values in a length-3 vector for each image. This vector represents the image as $-1$, 00, or $1$, based on the comparison with the thresholds.
@@
A neural network with a hidden layer of depth 1, width 2, and the Sigmoid activation function $\sigma(x)=\frac{1}{1+e^{-x}}$ is used for classification. We approximate the sigmoid activation function with a piecewise constant function. The breakpoints in this experiment is $\{-8,-4,0,4,8\}$ and the value of the piecewise linear function in each interval is equal to the function value of sigmoid function at the midpoint of that interval.
Appendix F Principles of STE and BinaryConnect
Straight-Through Estimator (STE) [47, 55]
The Straight-Through Estimator (STE) enables gradient-based training for non-differentiable operations, such as quantization, in neural networks. During the forward pass, a weight $x_{i}$ is quantized as:
$\hat{x}_{i}=s\cdot\text{clamp}\left(\left\lfloor\frac{x_{i}}{s}\right\rfloor,n,p\right),$ (99)
[…]
$\frac{\partial[\omega]}{\partial\omega}=1.$ (100)
This allows the gradient to “pass through” the non-differentiable operation, enabling backpropagation for quantized neural networks.
BinaryConnect [48]
BinaryConnect trains neural networks with binary weights, to reduce computational and memory costs. It maintains real-valued weights $w$ during training, which are binarized during the forward and backward passes using the sign function:
$w_{b}=\text{sign}(w),$ (101)
where $w_{b}\in\{-1,+1\}$. In the forward pass, binary weights $w_{b}$ are used to compute activations. In the backward pass, the gradient of the loss $C$ with respect to the binary weights, $\frac{\partial C}{\partial w_{b}}$, is computed and used to update the real-valued weights. This allows BinaryConnect to accumulate gradient updates in the real-valued weights while operating with binary weights.
Figure 6: Schematic of the coherent Ising machine (CIM) structure, utilizing optical parametric oscillators and phase-sensitive amplification to solve optimization problems.
Coherent Ising Machine

### v4

@@ Appendix B Error Bound of Piecewise Linear Approximation in FNN
To determine $n$, assume $\sigma$ has $\|\sigma^{\prime\prime}\|_{\infty}\leq M$. Construct $\hat{\sigma}$ over [R,R][-R,R] with $n$ equal segments, each of width $h=\frac{2R}{n}$. The interpolation error is:
@@
Appendix C Spline Quantumization Protocol
Quantum computing faces the dual challenge of optimizing not only QUBO models but also highly nonlinear optimization problems that are prevalent in scientific and industrial applications. The significance of this challenge lies in the ubiquity of problems featuring intricate, non-linear relationships among variables in real-world scenarios. This fact underscores the urgency for quantum algorithms capable of navigating and optimizing complex, highly nonlinear landscapes.
Formally, we consider the following optimization problem:
@@ Appendix C Spline Quantumization Protocol
where $f(\cdot)$ is a function of arbitrary form. This formulation expands the potential for quantum computing to address real-world optimization challenges. A powerful method for approximating highly nonlinear functions is spline interpolation. Splines are piecewise linear functions that provide a smooth and flexible fit to data points, thereby transforming complex optimization landscapes into more tractable forms suitable for quantum optimization algorithms. The spline fitting process involves the following steps:
@@
C.1 Piecewise Constant Segment
@@
We have now derived a QUBO model capable of approximately solving optimization problems with arbitrary objective functions in (C). However, it is important to note that in representing $\beta_{i}$, we introduced two inequality constraints. Here, we introduce the following two approaches:
@@
    It is important to note that the higher-order terms cancel out when the expression is expanded.
@@
Building upon the versatility of spline approximation in simplifying complex optimization landscapes, we use the following sign function as an example to illustrate the efficiency of our approach.
@@
$\text{sign}(x)$ (11 if $x\geq 0$ else 1-1) can be represented as
@@
Compared with the model in ref.[76], we do not need to introduce the intermediate variable. Additionally, when representing the sign activation function, we do not need to introduce auxiliary variables for order reduction [76]. These optimizations can save $O(WDN\log W)$ bits, where $W,D,N$ denote the network width, network depth, and dataset size, respectively. Indeed, we can reduce the coefficients of the highest-order terms in the model’s qubit count by $50\%$.
Appendix D Details of Convex Formulation
We consider a fixed dataset $\mathcal{S}=\{(\mathbf{x}^{(s)},y^{(s)})\}_{s=1}^{N}$ and an $L$-layer feedforward network with widths $m_{0},\dots,m_{L}$.
McCormick linear inequalities.
We now write the four inequalities in the global form $\mathbf{N}^{(s,\ell)}_{\mathrm{Mc}}\mathbf{u}\leq\mathbf{n}^{(s,\ell)}_{\mathrm{Mc}},$ with explicit block matrices as following.
For fixed $(s,\ell)$, we define
$\bm{\delta}^{(\ell)}_{W}\in\{0,1\}^{m_{\ell}m_{\ell-1}R_{\ell}},\quad\mathbf{a}^{(s,\ell-1)}\in\mathbb{R}^{m_{\ell-1}},\quad\mathbf{v}^{(s,\ell)}\in\mathbb{R}^{m_{\ell}m_{\ell-1}R_{\ell}},$ (106)
where $\mathbf{v}^{(s,\ell)}$ stacks $v^{(s,\ell)}_{jk,r}$ over $(j,k,r)$ in a consistent order, and $\bm{\delta}^{(\ell)}_{W}$ stacks the matching bits $\delta^{(\ell)}_{jk,r}$ in the same order. Define the Kronecker matrices
$\underline{\mathbf{A}}^{(s,\ell-1)}:=\mathbf{I}_{m_{\ell}}\,\otimes\,\mathrm{Diag}\!\big(\underline{\mathbf{a}}^{(s,\ell-1)}\big)\,\otimes\,\mathbf{I}_{R_{\ell}},\quad\overline{\mathbf{A}}^{(s,\ell-1)}:=\mathbf{I}_{m_{\ell}}\,\otimes\,\mathrm{Diag}\!\big(\overline{\mathbf{a}}^{(s,\ell-1)}\big)\,\otimes\,\mathbf{I}_{R_{\ell}},$ (107)
and
$\mathbf{P}_{a}:=\mathbf{I}_{m_{\ell}}\,\otimes\,\big(\mathbf{1}_{R_{\ell}}^{\top}\otimes\mathbf{I}_{m_{\ell-1}}\big)\ \in\ \mathbb{R}^{(m_{\ell}m_{\ell-1}R_{\ell})\times m_{\ell-1}}.$ (108)
For each row indexed by $(j,k,r)$, $\mathbf{P}_{a}$ selects $a^{(s,\ell-1)}_{k}$ and replicates it across bits.
Using these blocks, the four inequalities can be written compactly as one stacked system
$\mathbf{N}^{(s,\ell)}_{\mathrm{Mc}}\begin{bmatrix}\bm{\delta}^{(\ell)}_{W}\\ \mathbf{a}^{(s,\ell-1)}\\ \mathbf{v}^{(s,\ell)}\end{bmatrix}\;\leq\;\mathbf{n}^{(s,\ell)}_{\mathrm{Mc}},$ (109)
where
$\mathbf{N}^{(s,\ell)}_{\mathrm{Mc}}=\begin{bmatrix}-\overline{\mathbf{A}}^{(s,\ell-1)}&\mathbf{0}&\ \ \mathbf{I}\\ \underline{\mathbf{A}}^{(s,\ell-1)}&\mathbf{0}&-\mathbf{I}\\ -\underline{\mathbf{A}}^{(s,\ell-1)}&-\mathbf{P}_{a}&\ \ \mathbf{I}\\ \ \ \overline{\mathbf{A}}^{(s,\ell-1)}&\ \ \mathbf{P}_{a}&-\mathbf{I}\end{bmatrix},\qquad\mathbf{I}:=\mathbf{I}_{m_{\ell}m_{\ell-1}R_{\ell}}.$ (110)
The right-hand side vector $\mathbf{n}^{(s,\ell)}_{\mathrm{Mc}}$ stacks the corresponding bound terms:
$\mathbf{n}^{(s,\ell)}_{\mathrm{Mc}}=\begin{bmatrix}\mathbf{0}\\ \mathbf{0}\\ -\underline{\mathbf{a}}^{(s,\ell-1)}_{\mathrm{rep}}\\ \ \ \overline{\mathbf{a}}^{(s,\ell-1)}_{\mathrm{rep}}\end{bmatrix}$, (111)
where
$\underline{\mathbf{a}}^{(s,\ell-1)}_{\mathrm{rep}}:=\Big(\mathbf{I}_{m_{\ell}}\otimes(\mathbf{1}_{R_{\ell}}\otimes\mathbf{I}_{m_{\ell-1}})\Big)\,\underline{\mathbf{a}}^{(s,\ell-1)},\qquad\overline{\mathbf{a}}^{(s,\ell-1)}_{\mathrm{rep}}:=\Big(\mathbf{I}_{m_{\ell}}\otimes(\mathbf{1}_{R_{\ell}}\otimes\mathbf{I}_{m_{\ell-1}})\Big)\,\overline{\mathbf{a}}^{(s,\ell-1)}.$ (112)
Each block row corresponds exactly to one of the four scalar McCormick envelopes,
  • Row block 1 enforces $v^{(s,\ell)}_{jk,r}-\overline{a}^{(s,\ell-1)}_{k}\delta^{(\ell)}_{jk,r}\leq 0$.
  • Row block 2 enforces $v^{(s,\ell)}_{jk,r}+\underline{a}^{(s,\ell-1)}_{k}\delta^{(\ell)}_{jk,r}\leq 0$.
  • Row block 3 enforces $v^{(s,\ell)}_{jk,r}-a^{(s,\ell-1)}_{k}-\underline{a}^{(s,\ell-1)}_{k}\delta^{(\ell)}_{jk,r}\leq-\underline{a}^{(s,\ell-1)}_{k}$, which appears as $\underline{\mathbf{A}}^{(s,\ell-1)},\ -\mathbf{P}_{a},\ \mathbf{I}]\leq-\underline{\mathbf{a}}^{(s,\ell-1)}_{\mathrm{rep}}$.
  • Row block 4 enforces $v^{(s,\ell)}_{jk,r}+a^{(s,\ell-1)}_{k}+\overline{a}^{(s,\ell-1)}_{k}\delta^{(\ell)}_{jk,r}\leq\overline{a}^{(s,\ell-1)}_{k}$, which appears as $\overline{\mathbf{A}}^{(s,\ell-1)},\ \mathbf{P}_{a},\ -\mathbf{I}]\leq\overline{\mathbf{a}}^{(s,\ell-1)}_{\mathrm{rep}}$.
Reconstructing the affine layer.
Let $\mathbf{C}^{(\ell)}_{\mathrm{bit}}\in\mathbb{R}^{m_{\ell}m_{\ell-1}\times m_{\ell}m_{\ell-1}R_{\ell}}$ place the bit coefficient $c^{(\ell)}_{jk,r}$ at position $(j,k,r)$, and define
$\mathbf{u}^{(s,\ell)}_{\!\mathrm{bil}}\ =\ \mathbf{C}^{(\ell)}_{\mathrm{bit}}\mathbf{v}^{(s,\ell)}+\big(\mathbf{I}_{m_{\ell}}\otimes\mathrm{Diag}(\mathbf{a}^{(s,\ell-1)})\big)\mathbf{d}^{(\ell)}_{W},\qquad\mathbf{z}^{(s,\ell)}\ =\ \underbrace{\big(\mathbf{I}_{m_{\ell}}\otimes\mathbf{1}_{m_{\ell-1}}^{\top}\big)}_{\mathbf{S}^{(\ell)}_{\mathrm{sum}}}\mathbf{u}^{(s,\ell)}_{\!\mathrm{bil}}+\mathbf{b}^{(\ell)}.$ (113)
All relations in (113) are linear. The corresponding rows are assembled into a block-sparse matrix $\mathbf{M}_{\mathrm{aff}}$ acting on $\mathbf{u}$.
Activation and output selection.
Fix a grid $M_{0}<\cdots<M_{n}$, we define $\mathbf{M}_{L}=[M_{0},\dots,M_{n-1}]$, $\mathbf{M}_{U}=[M_{1},\dots,M_{n}]$, and $\mathbf{t}=[(\sigma(M_{i-1})+\sigma(M_{i}))/2]_{i}$. For each neuron $(s,\ell,j)$, we have
  • One-hot normalization:
    [ $\underbrace{\big[\ \mathbf{1}^{\top}\ \big]}_{\mathbf{m}^{\top}_{\text{1hot}}}\,\bm{\beta}^{(s,\ell)}_{z,j}\;=\;\underbrace{1}_{m_{\text{1hot}}}.$ (114)
    This row contributes to the global equality system $\mathbf{M}_{q}\mathbf{u}=\mathbf{m}_{q}$ by selecting the block $\bm{\beta}^{(s,\ell)}_{z,j}$ in $\mathbf{u}$ with coefficient vector $\mathbf{1}^{\top}$, and zero elsewhere.
  • Activation value tying:
    $\underbrace{\big[\ -(\mathbf{t}^{\top})\ \ \ 1\ \big]}_{\mathbf{m}^{\top}_{\text{act}}}\,\begin{bmatrix}\bm{\beta}^{(s,\ell)}_{z,j}\\ a^{(s,\ell)}_{j}\end{bmatrix}\;=\;0.$ (115)
    Equivalently $a^{(s,\ell)}_{j}-\mathbf{t}^{\top}\bm{\beta}^{(s,\ell)}_{z,j}=0$. This also enters $\mathbf{M}_{q}\mathbf{u}=\mathbf{m}_{q}$ with right-hand side 0.
  • Lower and upper interval bound:
    $\underbrace{\big[\ -\mathbf{M}_{L}\ \ \ 1\ \big]}_{\mathbf{n}^{\top}_{\text{low}}}\,\begin{bmatrix}\bm{\beta}^{(s,\ell)}_{z,j}\\ z^{(s,\ell)}_{j}\end{bmatrix}\;\geq\;0,\quad\text{i.e.}\quad z^{(s,\ell)}_{j}-\mathbf{M}_{L}\bm{\beta}^{(s,\ell)}_{z,j}\geq 0.$ (116)
    $\underbrace{\big[\ \mathbf{M}_{U}\ \ \ -1\ \big]}_{\mathbf{n}^{\top}_{\text{up}}}\,\begin{bmatrix}\bm{\beta}^{(s,\ell)}_{z,j}\\ z^{(s,\ell)}_{j}\end{bmatrix}\;\geq\;0,\quad\text{i.e.}\quad\mathbf{M}_{U}\bm{\beta}^{(s,\ell)}_{z,j}-z^{(s,\ell)}_{j}\geq 0.$ (117)
    These two rows together form the neuron-local block of the global inequality system $\mathbf{N}_{r}\mathbf{u}\leq\mathbf{n}_{r}$. Each such inequality will later receive a nonnegative slack variable $s_{r}$ so that it can be rewritten as an equality $\mathbf{N}_{r}\mathbf{u}+s_{r}=\mathbf{n}_{r}$ before CP lifting.
For the network output $a^{(s,L)}$, we introduce another one-hot vector $\bm{\beta}^{(s)}_{a}$ and apply the same construction, and the corresponding rows contribute additional blocks in the convex formulation in the same way.
@@ Appendix D Details of Convex Formulation
Appendix E Proof of Convergence of QCGD with Random Quantum Oracle
@@
Proof: Similar to ref.[49], the algorithm utilizes an augmented Lagrangian function, defined as
@@
$\mathcal{Q}_{\alpha}(\mathbf{V},\mathbf{z})=\mathrm{Tr}(\mathbf{C}\mathbf{V})+\mathbf{z}^{\top}(\mathcal{L}\mathbf{V}-\mathbf{v})+\frac{\alpha}{2}\|\mathcal{L}\mathbf{V}-\mathbf{v}\|^{2}\quad\text{for}\penalty 10000\ \mathbf{V}\in\Delta^{p},$ (118)
@@
The QCGD algorithm begins by initializing $\alpha$, $\mathbf{V}$, and $\mathbf{z}$, and then proceeds iteratively through alternating primal and dual updates. In the primal step, the algorithm computes the gradient of the augmented Lagrangian with respect to $\mathbf{V}$ and identifies a direction that minimizes the linearized loss, which involves solving a standard QUBO problem to determine the update direction. The primal variable $\mathbf{V}$ is subsequently adjusted by moving along this direction. In the dual step, the gradient of the augmented Lagrangian with respect to $\mathbf{z}$ is computed, and $\mathbf{z}$ is updated using gradient ascent. Throughout the iterations, the penalty parameter $\alpha$ is increased as $\alpha_{i}=\alpha_{0}\sqrt{\delta i+1}$ to ensure that $\mathbf{V}$ converges to a feasible solution that satisfies the constraints.
@@
$\text{Objective-gap}_{T}=\mathrm{Tr}(\mathbf{C}\mathbf{V}_{T+1})-\mathrm{Tr}(\mathbf{C}\mathbf{V}_{\star})=O\Big(\frac{(1+\varepsilon)}{\delta^{3/2}\sqrt{T}}\Big)$ (137)
@@
Appendix F Additional Experimental Results
F.1 Additional Details of the Fashion MNIST Binary Classification Results
@@
The training set consists of $12000$ images ($6000$ coat and $6,000$ sandal images), and the test set consists of $2000$ images ($1000$ coat and $1000$ sandal images). Each image is divided into $3$ columns vertically, resulting in two $28\times 9$ grids and one $28\times 10$ grid. We then calculate the number of zero pixels in each column and use two threshold values to determine the values in a length-3 vector for each image. This vector represents the image as 1-1, 0, or $1$, based on the comparison with the thresholds.
@@
A neural network with a hidden layer of depth 1, width 2, and the Sigmoid activation function $\sigma(x)=\frac{1}{1+e^{-x}}$ is used for classification. We approximate the sigmoid activation function with a piecewise constant function. The breakpoints in this experiment are $\{-8,-4,0,4,8\}$, and the value of the piecewise linear function in each interval is equal to the function value of the sigmoid function at the midpoint of that interval.
F.2 Robustness of hybrid approach: impact of coefficient matrix rounding
To investigate the convergence behavior of the hybrid algorithm in neural network training under different matrix rounding precision levels, we systematically rounded the matrix coefficients to digit precisions ranging from 1 to 10 decimal places. The algorithm was then executed for each precision setting until convergence to the optimal solution was achieved. Specifically, matrix elements were rounded to 1 through 10 significant digits.
Figure 6(a) displays the QUBO matrix evolution using circular heatmaps, highlighting the impact of precision levels on iterative trajectories. Despite the differences revealed in these trajectories, our results confirm that all trajectories converge to the optimal solution, consistent with the theoretical convergence of the QCGD algorithm. Notably, even under minimal precision conditions (e.g., 1-digit rounding), matrix adjustments in later iterations remain marginal. These consistently small perturbations across all precision levels suggest that a lazy evaluation strategy for QCGD could be viable. Such an approach would minimize redundant quantum circuit executions, thereby improving computational efficiency on quantum hardware.
The smoothed absolute differences of the objective value and constraint residual relative to the 10-digit precision baseline were recorded, with results visualized in Figure 6(b) and 6(c). This setup allowed us to assess the impact of numerical precision on convergence trajectories. Both subplots demonstrate that differences across all precision levels remain small, typically less than $10^{-2}$, indicating the hybrid algorithm’s robustness to precision variations. A downward trend is observed in the difference of constraint residual, with differences decreasing sharply after initial fluctuations and stabilizing post-1500 iterations, confirming convergence to a consistent optimal solution regardless of precision, as supported by prior results. The rapid initial decline in the constraint residual difference, particularly for lower precisions (e.g., 1-3 digits), suggests early sensitivity to numerical accuracy, while the differences remain stably small in later iterations. The near-overlap of higher precision curves (e.g., 7-9 digits) highlights diminishing returns from increased precision.
F.3 Analysis of Randomized Rounding Procedure
Let $\bar{x}\in[0,1]^{n}$ be the DLBO solution and denote by $F_{\mathrm{DLBO}}$ its objective value. Consider $K$ independent randomized roundings of $\bar{x}$,
$X^{(1)},\dots,X^{(K)},\qquad X^{(k)}_{j}\sim\mathrm{Bernoulli}(\bar{x}_{j})\ \text{independently for }j=1,\dots,n,$
and write $F^{(k)}:=F(X^{(k)})$ as well as $F_{\min}:=\min_{1\leq k\leq K}F^{(k)}$.
As the objective $F$ is defined as the sum of per–sample losses measuring the discrepancy between the predicted labels and the true labels, we claim that $F$ is bounded and thus satisfies the bounded–difference property, i.e., there exist constants $\Delta_{1},\dots,\Delta_{n}$ such that changing only the $j$-th coordinate of the argument changes $F$ by at most $\Delta_{j}$. McDiarmid’s inequality [77] then yields, for any $t>0$ and each fixed $k$,
$\Pr\!\left(F^{(k)}\geq\mathbb{E}[F(X)]+t\right)\;\leq\;\exp\!\left(-\frac{2t^{2}}{\sum_{j=1}^{n}\Delta_{j}^{2}}\right).$
Using the independence of the $K$ samples, we obtain
$\Pr\!\left(F_{\min}\geq\mathbb{E}[F(X)]+t\right)=\Pr\!\left(\bigcap_{k=1}^{K}\bigl\{F^{(k)}\geq\mathbb{E}[F(X)]+t\bigr\}\right)\leq\left[\exp\!\left(-\frac{2t^{2}}{\sum_{j=1}^{n}\Delta_{j}^{2}}\right)\right]^{K}=\exp\!\left(-\frac{2Kt^{2}}{\sum_{j=1}^{n}\Delta_{j}^{2}}\right).$
Setting the right-hand side equal to $\delta\in(0,1)$ and solving for $t$ gives
$F_{\min}\;\leq\;\mathbb{E}[F(X)]+\sqrt{\frac{1}{2K}\Bigl(\sum_{j=1}^{n}\Delta_{j}^{2}\Bigr)\log\frac{1}{\delta}}\quad\text{with probability at least }1-\delta.$
We can assume that the discrete objective $F$ is a quadratic polynomial with zero diagonal, i.e.,
$F(x)\;=\;x^{\top}Qx+c^{\top}x,\qquad x\in\{0,1\}^{n},$
where $Q\in\mathbb{R}^{n\times n}$ is symmetric with $Q_{ii}=0$ for all $i$ and $c\in\mathbb{R}^{n}$. Because for binary variables $x_{i}^{2}=x_{i}$, so each diagonal coefficient $Q_{ii}$ can be absorbed into the linear term. Let $\bar{x}\in[0,1]^{n}$ be the relaxed solution and consider the independent Bernoulli rounding
$X_{j}\sim\mathrm{Bernoulli}(\bar{x}_{j}),\quad j=1,\dots,n,$
with all coordinates independent. Then $\mathbb{E}[X_{j}]=\bar{x}_{j}$ and, for $i\neq j$,
$\mathbb{E}[X_{i}X_{j}]\;=\;\mathbb{E}[X_{i}]\mathbb{E}[X_{j}]\;=\;\bar{x}_{i}\bar{x}_{j}.$
Using these identities, we can compute
$\mathbb{E}[F(X)]=\mathbb{E}[X^{\top}QX]+\mathbb{E}[c^{\top}X]=\sum_{i,j}Q_{ij}\,\mathbb{E}[X_{i}X_{j}]+\sum_{j}c_{j}\,\mathbb{E}[X_{j}].$
Because $Q_{ii}=0$, the quadratic term reduces to
$\sum_{i\neq j}Q_{ij}\,\mathbb{E}[X_{i}X_{j}]=\sum_{i\neq j}Q_{ij}\,\bar{x}_{i}\bar{x}_{j}=\bar{x}^{\top}Q\bar{x},$
and the linear term is
$\sum_{j}c_{j}\,\mathbb{E}[X_{j}]=\sum_{j}c_{j}\bar{x}_{j}=c^{\top}\bar{x}.$
Therefore,
$\mathbb{E}[F(X)]=\bar{x}^{\top}Q\bar{x}+c^{\top}\bar{x}=F(\bar{x}).$
Substituting this bound into the previous display yields the comparison between the best-rounded solution and the DLBO objective value, i.e.,
$F_{\min}\;\leq\;F(\bar{x})+\sqrt{\frac{1}{2K}\Bigl(\sum_{j=1}^{n}\Delta_{j}^{2}\Bigr)\log\frac{1}{\delta}}$
with probability at least $1-\delta$.
Figure 7: Schematic of the setup of the coherent Ising machine (CIM).
Appendix G Principles of Coherent Ising Machine